OpenStudy (anonymous):

what is the slope of a line that is perpendicular to 6x+3y=-9?

OpenStudy (ash2326):

first find the slope of this line 3y=-9-6x y=-2x-3 y=mx+c m=slope c=intercept on y axis m (slope) =-2 let the line perpendicular to this have a slope of a so a *(-2)=-1 a=1/2 slope of line perpendicular to this line =1/2 or 0.5

OpenStudy (ash2326):

if two lines are perpendicular then the product of their slopes is -1

OpenStudy (anonymous):

want to help me with rest?

OpenStudy (ash2326):

yeah sure

OpenStudy (anonymous):

what is the equation of the line that passes through the points (1,2/3) and (-1,-1)?